On January 1, 2018, ABC & Co. issues convertible bonds with a maturity of 5 years. The par value of the bonds is $400,000, the coupon rate is 6%, and the compounding period is semi-annual with interest paid on June 30th and December 31st. The market prices these bonds using an interest rate (effective rate) of 4% compounded semi-annually. Each $1,000 bond is convertible to 100 shares of ABC & Co. common stock.

1. On July 1, 2018, the company has excess cash and buys back one-half of the bonds. The market interest rate on July 1, 2018 is 8% compounded semi-annually. Prepare the journal entry for the early extinguishment of one-half of the bonds on July 1, 2018.

2.On January 1, 2019, the bondholders exercised their conversion option on the one-half of the bonds still outstanding. The market price of ABC & Co. shares is $12. Prepare the journal entry for the conversion of one-half of the bonds on January 1, 2019 both the book value method and the market value method.
Answer #1
1. Journal Entry for Early Extinguishment of one-half of the bonds.
Date Particulars Debit ($) Credit ($)
July 1, 2018 Bonds Payable $200,000
Premium on Bonds Payable (Expn. 1) $16,325
       Cash ($ 370,260*1/2) $185,130
       Gain on Early Extinguishment $31,195
Fair Value of Bonds as on July 1, 2018
Fair Value = PVA(n=9,r=4%)*400000*6%*6/12+PV(n=9,r=4%)*400000
Fair Value = $ 89,224 + $ 281,035
Fair Value = $ 370,260
Expn. 1 Premium on Bonds Payable to be written off
Fair Value(on Jan.1,2018)= PVA(n=10,r=2%)*400000*6%*6/12+PV(n=10,r=2%)*400000
Fair Value(on Jan.1,2018)= $ 107,791 + $ 328,139
Fair Value(on Jan.1,2018)= $ 435,930
Therefore, Premium on Bonds Payable = $ 435,930 - $ 400,000 = $ 35,930
Now, Book Value of Bonds as on Jan. 1, 2018
Book Value = $ 435,930 + (($435,930*2%)-$12000)
Book Value = $ 435,930 - $ 3280
Book Value = $ 432,650
Premium on Bonds Payable to be written off
        = ($ 432,650 - $ 400,000)/2
        = $ 16,325
2. Journal Entry on Conversion of the remaining Half of the bonds.
Book Value Method
Date Particulars Debit ($) Credit ($)
Jan 1, 2019 Bonds Payable $200,000
Premium on Bonds Payable (Expn. 2) $14,625
       Common Stock $200,000
       Gain on Conversion $14,625
Market Value Method
Date Particulars Debit ($) Credit ($)
Jan 1, 2019 Bonds Payable $200,000
Premium on Bonds Payable (Expn. 2) $14,625
Loss on Conversion $25,375
       Common Stock $240,000
Expn. 2 Premium on Bonds Payable to be written off
Fair Value(on Jan.1,2019)= PVA(n=8,r=4%)*200000*6%*6/12+PV(n=8,r=4%)*200000
Fair Value(on Jan.1,2019)= $ 46,397 + $ 146,138
Fair Value(on Jan.1,2019)= $ 192,535
Now, Book Value of Bonds as on Jan. 1, 2019
Book Value = $ 435,930 + (($435,930*2%)-$12000) -$ 216,325 + (($192,535*4%)-$6000)
Book Value = $ 435,930 - $ 3,280 - $ 216,325 - $ 1,702
Book Value = $ 214625
Premium on Bonds Payable to be written off
        = $ 214,625 - $ 200,000
        = $ 14,625